ID |
Date |
Author |
Type |
Category |
Subject |
2477
|
Tue Jan 5 10:26:32 2010 |
Alberto, Steve | Omnistructure | Environment | Added new wall cable-racks |
we hung two new WALL cable racks. One is on the pillar next to the Sp table, the other is next to the PSL computer rack.
To do that we had to drill holes in the wall since the simple screws weren't strong enough to keep them up.
One of the racks, the yellow, is dedicated to 4-pin lemos and other thick cables.
 
|
2476
|
Tue Jan 5 09:18:38 2010 |
Alberto | Omnistructure | Electronics | Universal PDH Box Stored in the RF Cabinet |
FYI: I stored the Universal PDH boxes in the RF cabiner in the Y arm. |
2475
|
Tue Jan 5 01:31:09 2010 |
Jenne | Update | WienerFiltering | New Wiener Filters installed in PEM IIR matrix on OAF screen |
Using the techniques employed at LLO, and then by Rana here at the 40m a few weeks ago, Wiener filters have been installed on the inputs of all of the PEM IIR channels which are hooked up to the 110B PEM ADCU. Some slight modifications have taken place to the code, and it's all been checked in to the 40m svn.
I have installed the filters into: All 6 Wilcoxon accelerometers, the Ranger seismometer, and one of the Guralps (GUR1). The other Guralp is currently connected through the ASS/OAF machine's ADC, so it's not used in this test. The filters are all labeled "Wiener", and are FM1 in the C1:ASS-TOP_PEMIIR_## filter banks.
The first figure below is the output of the Wiener Filter calculation program. It shows the uncorrected MC_L (black) and the corrected MC_L (red), using the optimal wiener filter. This is as good as we should be able to do with these sensors in these positions.
The second figure is a DTT shot of me trying out the nifty new filters. They seem to maybe do a teensy bit on the microseism, but otherwise it's a bit unremarkable. Hopefully I'll get better subtraction during the day, when the base level for MC_L is higher. Here, Black is uncorrected MC_L, Red is the corrected MC_L, Blue is the actuation channel, and green is an example seismometer channel to illustrate the ground motion at the time.
For posterity, since it's not all in one elog that I can find, the order of operations to install a Wiener Feed Forward filter is as follows:
1. (When you can borrow the IFO) Take a very careful TF of the plant, between your actuation point and your error signal readout. At the 40m, this means between C1:ASS-TOP_SUS_MC1_EXC and C1:IOO-MC_L, since we actuate by pushing on the MC1 coils. At the sites / future 40m, this would be between the HEPI (or STACIS) and the error signal. The limit of how good your Feed Forward can do will depend heavily on how good this TF is. Coherence should be above ~0.95 for all points. Export this data from DTT as a .txt file, using units "Complex (abs/rad)".
2. Run fitMC12MCL.m (or equivalent wrapper file) to fit the transfer function you just took with some Poles and Zeros. Make sure to edit the wrapper file with your new .txt file's name so you're getting a fresh TF (if you've just taken one).
3. (Again, when you can borrow the IFO) Run getMCdata.m (or equivalent) to fetch witness channel data and error signal data. At the 40m, this usually means C1:IOO-MC_L, and witness sensors which are around the MC chambers. This data should be taken at a time when the cavity is locked, but pretty much on it's own. (i.e. probably shouldn't have Common Mode feedback on the MC - so the MC should be locked, but not the full IFO, for example.)
4. Run c1winoiir.m (the main program here, which contains some of these notes). This will take in the TF data you've fit, and the witness channel data you have, and calculate the optimal combination of Wiener filters for your witness channels. It pre-filters your witness data by your TF, then calculates the Wiener filter. The resulting FIR filters are saved in a file.
5. Run firfit.m This will take the FIR Wiener filters you've just created, and convert them conveniently to IIR filters, in a format to be copied directly into Foton. For each witness channel, you'll get the IIR filter in 2 formats: the first is for copying into the Foton .txt file (ex ASS.txt), and the second is for copying into the Foton gui, in the "Command" box on the filter design screen. The "o" at the end of the copy-able filter indicates to Foton that it is a Z-Plane Online filter. Copy filters appropriately (there should be a line preceding each set of SOS filter formats to indicate which channel this Wiener filter is for...these channel names are extracted from your getMCdata.m)
6. Save your Foton file, and update your Coefficients in MEDM. Enable your outputs to actuators, and watch magic happen!
On the To-Do list:
Check the transfer of signal btwn PEMIIR matrix and the 9x1 'matrix' that sends the signals to the SUS inputs. In the SimuLink, the input to the 9x1 matrix is a bundle of 8 numbers (the 8 outputs of the PEMIIR matrix), but it looks like it only pays attention to the first one. Need to figure out how to make it realize that it's a bundle, not a single number.
Also, the OAF up / down scripts don't seem to be working on any of the control room computers. This needs to be checked in to / fixed (but not tonight....) |
Attachment 1: MCwino-FFtest_4Jan10_sameFiltersAsinEPICS.png
|
|
Attachment 2: OAF-FF_test_4Jan10.png
|
|
2474
|
Mon Jan 4 17:26:01 2010 |
Mott | Update | General | T & R plots for Y1 and Y1S mirrors |
The most up-to-date T and R plots for the Y1 and Y1S mirrors, as well as a T measurement for the ETM, can be found on:
http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/Optics/RTmeasurement
|
2473
|
Mon Jan 4 17:21:30 2010 |
Jenne | Configuration | IOO | Elusive Mode Matching Solution found! |
I think I have finally found a Mode Matching solution for our new Input Mode Matching Telescope! And after looking at the layout diagram with Koji and Raffaele, it seems like all of the optics will fit into the chambers / onto the tables (not true as of last week).
3. RoCMMT1 is -5m
RoCMMT2 is 8m,
with the MMTs 1.89m apart.
This is a 1.6x telescope.
MMT2 is 2.2641m from the PRM
MMT1 is 2m from MC3.
The Condition Number for this optical chain is 89219047.5781.
This layout is very similar to the one that Koji posted on the wiki yesterday: Upgrade09/Optical Layout. The difference is that I want to move MMT1 ~20cm closer to the MC13 table, so just on the other side of the main red beam that goes directly to PRM. There is plenty of space there, so it should be all good. The tricky bit is that the flat steering mirrors fit into things now while they are piezos, but they will be trickier to fit if we make them into Tip Tilts. But I have full faith in Koji's amazing optical table layout skills, that he can make it happen.
Unless there are major objections, I think this is the MMT that we're going to go with. (So speak now or forever hold your peace.) The angle between tilt and translation isn't quite what we'd like it to be (at ~18deg), but it's not too terrible. And we still have 99.5% overlap which is very important. |
Attachment 1: Awesome_MM_Solution.png
|
|
2472
|
Mon Jan 4 09:52:40 2010 |
rana | Configuration | Cameras | ITMX camera and PSL channels |
I fixed up the ITMX camera like we did for ITMY recently (removed T's and added strain relief - the lens was already OK).
I also updated the .SCAN field for the RMTEMP and RCTEMP channels to 0.1 second. This had been done via probe but was wiped out after reboot previously, because I forgot to update the psl.db file. |
2471
|
Sun Jan 3 08:23:39 2010 |
rana | Configuration | CDS | autoburt.pl 'fixed' for post 2009 years |
Tobin & Keith pointed out in the LLO ilog that there was a code bug in the autoburt.pl script for autoburts.
I edited the autoburt.pl script so that it will work from now until 2099 (by which time we may no longer be using this version of perl):
nodus:autoburt>diff autoburt.pl~ autoburt.pl
234c234
< $thisyear = "200".$timestamp[5];
---
> $thisyear = "20".$timestamp[5];
The autoburt has not been working ever since 11PM on New Year's eve.
I ran it by hand and it seems to run fine. I noticed along the way that it was running on op340m (our old Sun Blade 150 machine). The autoburt.pl was pointing at /cvs/cds/bin/perl
which is Perl v5.0. I changed it to use '/usr/bin/env' and now points at '/usr/bin/perl' which is perl 5.8. It runs fine with the new perl:
op340m:scripts>time perl /cvs/cds/scripts/autoburt.pl >> /cvs/cds/caltech/logs/autoburtlog.log
5.37u 6.29s 2:13.41 8.7%
Also ran correctly, via cron, at 9AM.
|
2470
|
Wed Dec 30 22:17:07 2009 |
kiwamu | Update | General | Camera input and monitor output |
The input channels of the cameras and the output channels for the monitors are summarized on the wiki.
The channel table on the wiki is very helpful when you want to make a change in the video matrix.
thank you. |
2469
|
Wed Dec 30 20:33:36 2009 |
rana, alberto | Configuration | Cameras | ITMY & MC2 Camera work |
We restored the good state of the ITMY camera and neatened both the MC2 and ITMY camera.
The MC2 camera was driving a triple T jungle into some random cables and spoiling the image. We removed all T's and the MC2 camera now drives only The Matrix.
The ITMY camera was completely unmounted and T'd. So it was misaligned just by the force of gravity acting on its BNC cable. We swapped the lens for a reasonable sized one and remounted it in its can. We then used orange cable ties to secure the power and BNC cable for the MC2 and ITMY cameras so that tugging on the cables doesn't misalign the cameras. This is part of the camera's SOP.
No more driving 50 Ohm cables and T's with video cables, Steve! If you need a portable video, just use a spigot of the Matrix and then you can control it with a web browser.
  
I also wiped out the D40's memory card after uploading all of the semi-useful files to the Picasa page. |
2468
|
Wed Dec 30 18:01:03 2009 |
Alberto, Rana | Update | General | All watchdogs tripped this morning |
WQuote: |
This morning I found all the watchdogs had tripped during the night.
I restored them all.
I can't damp ITMX. I noticed that its driving matrix is all 1s and -1s as the the right values had been lost in some previous burtrestoring.
|
Rana fixed the problem. He found that the side damping was saturating. He lowered the gain a little for a while, waited for the the damping to slow down the optic and then he brought the gain back where it was.
He also upadted the MEDM screen snapshot. |
2467
|
Wed Dec 30 10:58:48 2009 |
Alberto | Update | General | All watchdogs tripped this morning |
This morning I found all the watchdogs had tripped during the night.
I restored them all.
I can't damp ITMX. I noticed that its driving matrix is all 1s and -1s as the the right values had been lost in some previous burtrestoring. |
2466
|
Wed Dec 30 09:57:53 2009 |
steve | Update | PSL | thirsty water chiller |
I added 600 cc of Arrowhead Distilled Water to the chiller.
60 days plot shows that about every ~ 10 days I have to add some.
Please check the water level yourself. |
Attachment 1: htemp60d.png
|
|
2465
|
Tue Dec 29 13:57:20 2009 |
Rana, Kiwamu, and Haixing | Update | Photos | Photos of video switch box |
Before we installed the video switch box, we also took some photos of it. We uploaded them onto the 40m Picasa.
Video Matrix
The first photo is the an entire view of the switch box. The following four photos are the details of the switch matrix.
The slideshow below is a dump of the last several months of photos from the Olympus. The originals have been deleted.
|
2464
|
Tue Dec 29 04:28:27 2009 |
kiwamu, rana, haixing | Update | Cameras | New Video Switch Installed |
We have installed the new Video Matrix.
Its still in an intermediate state, so don't try to "fix" anything before Kiwamu and I get back onto it in the afternoon.
The status so far is that we have removed the old switch (it was a 256 input x 128 output !! mux) and installed the new one in the same rack. We have hooked it up to the CDS network and have set up its matrix by using the web interface (i.e. NOT EPICS).
Along the way, we discovered that there is lack of impedance matching in the video all over the 40m. Video signals are RF and need to be treated that way. The PSL signals are T'd around and sent on 50 Ohm cables to high impedance monitor inputs.
We should eliminate any switches besides the new one (called Luciana) and control the PSL's Video Monitor from the main MUX interface. No more Rogue Video Switches.
Another couple of things we have found is about RCR camera.
(1) The long cable which connects the RCR camera box and the video matrix doesn't work. Although the signal is alive and we can see it by the local tv monitor nearby PSL.
(1) The reflected beam going to the camera is too weak to see in the monitor. We found a strange polarized cube splitter in front of the camera. We should modify it sooner or later. |
2462
|
Mon Dec 28 23:56:44 2009 |
kiwamu, rana | Update | Computers | add the HILO drift channels to the burt |
The HIGH and LOW channels are added into the burt request file "/target/c1losepics/autoBurt.req".
These values are used to colorize the alarm texts in the "C1DRIFT_MONITOR.adl" like a threshold. (the screenshot attached)
Hereafter these values will be automatically restored by the burt. Happy ! |
Attachment 1: Screenshot_DRIFTMON.png
|
|
2461
|
Mon Dec 28 18:35:27 2009 |
Alberto | Update | ABSL | Working on the AP table |
Quote: |
I opened the auxiliary laser's shutter.
I'm currently working on the AP table.
|
I finished working on the table.
I closed the AUX NPRO's shutter. |
2460
|
Mon Dec 28 15:34:14 2009 |
Alberto | Update | ABSL | Working on the AP table |
I opened the auxiliary laser's shutter.
I'm currently working on the AP table. |
2459
|
Mon Dec 28 15:19:19 2009 |
Alberto | Update | Computers | Burtrestored to Dec 26 at 20:00 |
Since it wasn't sure whether all the front-ends had been restored after the bootfest, I burtrestored everything to Dec 26 at 20:00.
Always keep in mind that to burtrestore c1dcuepics, the snapshot file has to be modified by hand by moving the last quote up to the line before the last. |
2458
|
Mon Dec 28 12:45:55 2009 |
Koji | Update | SUS | MC2 is having a bad day |
The MCL path of MC2 was in a strange state as the filters were activated as if it is in lock even though we had no lock. So I manually ran "mcdown". This reset the filters of the MCL path.
Quote: |
MC2 is having a bad day, and I'm not yet sure why. It's to do with the damping though. When the damping is off, after a little while it will settle to ~30mV or so on the Watchdog screen. When I enable all of the outputs and then turn on the damping, the optic gets kicked up. It's like there's a minus sign error somewhere, maybe in a bad burtrestore? This has been going on since I did my morning bootfest.
It's started to sit down and play nicely now. Is someone doing magic remotely that is fixing things that I hadn't figured out yet?
|
|
2457
|
Mon Dec 28 12:35:57 2009 |
Jenne | Update | SUS | MC2 is having a bad day |
MC2 is having a bad day, and I'm not yet sure why. It's to do with the damping though. When the damping is off, after a little while it will settle to ~30mV or so on the Watchdog screen. When I enable all of the outputs and then turn on the damping, the optic gets kicked up. It's like there's a minus sign error somewhere, maybe in a bad burtrestore? This has been going on since I did my morning bootfest.
It's started to sit down and play nicely now. Is someone doing magic remotely that is fixing things that I hadn't figured out yet? |
2456
|
Mon Dec 28 10:29:31 2009 |
Jenne | Update | Computers | Monday Morning Bootfest |
Nothing like a good ol' Bootfest to get back into the swing of things after vacation....
It was a regular bootfest, keying crates and running everyone's startup.cmd . There wasn't any RFM funny business which we had been dealing with a lot earlier in December (maybe Kiwamu took care of that part of things last night).
After finishing the bootfest, I tried to re-enable the watchdogs. I noticed that the optics weren't damping at all (not that any of them were swinging crazily, they just weren't damped like regular). This was traced to the OSEM sensor inputs and outputs being disabled on all of the suspensions' screens. I suspect that no burt-restoring happened after c1dcuepics was powercycled yesterday.
All of the optics are now happy as clams. |
2455
|
Mon Dec 28 01:17:01 2009 |
Koji | Update | Electronics | MCT QPD investigation |
Unfortunately, the signals for individual segments also suffer from the voltage drop as all of the low impedance amplifiers are hung from the same input.
In order to utilize the individual channels, we anyway have to remove the resistors for the VERT/HOR/SUM amps.
That is possible. But does it disable some fast channels for future ASC purposes?
Quote: |
This is indeed sad. But, we can perhaps bypass all of this by just using the individual segment outputs. According to the circuit diagram and the c1iool0 .db file, we should be able to just do the math on the segments and ignore the VERT/HOR/SUM signals completely. In that case, we can just use high impedance for the sum/diff buffers as Koji says and not suffer from the calibration errors at all I think.
|
|
2454
|
Sun Dec 27 23:44:59 2009 |
rana | Update | Electronics | MCT QPD investigation |
Quote: |
I found that MCT QPD has dependence of the total output on the position of the spot. Since the QPD needs the supply and bias voltages from the sum/diff amp, I could not separate the problems of the QPD iteself and the sum/diff amplifier by the investigation on Tuesday. On Wednesday, I investigated a generic quad photodiode interface module D990692.
|
This is indeed sad. But, we can perhaps bypass all of this by just using the individual segment outputs. According to the circuit diagram and the c1iool0 .db file, we should be able to just do the math on the segments and ignore the VERT/HOR/SUM signals completely. In that case, we can just use high impedance for the sum/diff buffers as Koji says and not suffer from the calibration errors at all I think. |
2453
|
Sun Dec 27 20:05:28 2009 |
kiwamu | Update | Computers | can not communicate with front-ends |
In this evening I found that fb40m has been down, then I restarted fm40m successfully.
However there still is a problem, the reflective memory can not communicate with some front-end CPUs ( such as c1iscey, c1susvme, ...etc.)
Right now I don't have any ideas about this, I am leaving them as they are now .... we can deal with them tomorrow.
The followers are what I did.
(1) ssh to fb40m then "pkill tpman"
(2) telnet to fb40m then typed "shutdown". ( These procedure are on the 40m wiki)
(3) make sure fb40m gets recovered while watching the medm screen C0DAQ_DETAIL.adl
(4) run the backup script in fb40m
(5) in order to fix the communication problem, physically turn off c1dcuepics and c0daqctrl
(6) keying some front-end CPUs. ---> still some of front ends indicate RED on the medm screen C0DAQ_DETAIL.adl ( figure attached )
|
Attachment 1: C0DAQ_DETAIL.png
|
|
2452
|
Sat Dec 26 19:22:13 2009 |
Koji | Summary | General | Mode coupling of two astigmatic beams |
In this note, amplitude and power couplings of two astigmatic (0,0)-th order gaussian modes are calculated. |
Attachment 1: mode_coupling.pdf
|
|
2451
|
Thu Dec 24 19:13:29 2009 |
Koji | Update | IOO | MCT QPD investigation |
I found that MCT QPD has a dependence of the total output on the position of the spot. Since the QPD needs the supply and bias voltages from the sum/diff amp, I could not separate the problems of the QPD itself and the sum/diff amplifier by the investigation on Tuesday. On Wednesday, I investigated a generic quad photodiode interface module D990692.
...I was so disappointed. This circuit was left uninvestigated and used so long time with the following sorrowful conditions.
- This circuit has 4 unbuffered inputs with input impedance of 300~400 Ohm. It's way too low!
- Moreover, those channels have different input impedances. Ahhhh.
- Even worse, the QPD circuit D990272 has output impedance of 50 Ohm.
- The PCB of this circuit has four layers. It is quite difficult to make modifications of the signal route.
- It is a headache: this circuit is "generic" and used in many places.
D990692 has 4 channel inputs that are not buffered. Each channel has two high impedance buffers but they are used only for the monitors. The signal paths have no buffer.
The differential amplifier is formed by R=1k Ohm. The inverted side of the input has 1kOhm impedance. The non-inverted side has 1.5kOhm impedance.
CH1: 10K // 1.5k // 1.5k // 1k = 411 Ohm
CH2: 10K // 1.5k // 1k // 1k = 361 Ohm
CH3: 10K // 1k // 1k // 1k = 323 Ohm
CH4: 10K // 1k // 1.5k // 1k = 361 Ohm
Considering the output impedance of 50Ohm for the QPD, those too low input impedances result in the following effect:
- Because of the voltage division, we suffer absolute errors of 10.8~13.4%. This is huge.
- Because of the input impedance differences, we suffer a relative error of 1.5%~3%. This is also huge.
Unfortunately, the circuit has no room to modify; the signal paths are embedded in the internal layer.
I decided to replace the resistors of the sum/diff amps from 1k to 10k. Also the input impedance of the buffer was removed as the input is terminated by the sum/diff amps in any case.This changes the input inpedance to the followings:
CH1: 15k // 15k // 10k = 4286 Ohm
CH2: 15k // 10k // 10k = 3750 Ohm
CH3: 10k // 10k // 10k = 3333 Ohm
CH4: 10K // 15k // 10k = 3750 Ohm
These yield the absolute error of 1.2-1.5%. The relative error is now 0.3%. I can accept these numbers, but later I should put additional terminating resistors to compensate the differencies.
So far I have modified the resistors for the MCT as the modification for a QPD needs 17 10k resistors.
Next thing I have to check is the dependence of the QPD outputs on the spot positions.
-----------------------------------------------
Edit: Feb 11, 2010
I talked with Frank and he pointed out that the impedances are not the matter but the gains of the each channels are the matters (after considering the output impedance of the QPD channels).
If we assume the ideal voltage sources at the QPD and the symmetric output impedances of 50Ohm, the gain of the each circuit are affected but the change should be symmetric.
He found that several things:
- The analog switch (MAX333) used in the QPD unit adds more output impedance (somewhat randomly!).
- The resistance of the sum/diff circuits may vary each other unless we use 0.1% resistors.
|
Attachment 1: D990692.png
|
|
2450
|
Thu Dec 24 01:25:29 2009 |
kiwamu | Update | Electronics | impedance analyzing |
The validation for high impedance measurement has been well done. 
The impedance measurement is one of the keys for designing the EOM circuit.
So far I was very struggling to measure the high impedance ( above several 1000 Ohm) at RF because the EOM circuit has a high impedance at its resonance.
Finally I realized that the measured impedance was suppressed by a parasitic resistance, which especially reduces the impedance at the resonance.
Also I found that we can extract the TRUE impedance data by subtracting the effect of the parasitic resistance from resultant data.
In order to confirm whether this subtraction works correctly or not, the impedance was directly re-measured with another analyzer for crosscheck.
The followers are details about the re-measurement.
(measurement )
The measurement has been performed with help from Peter and Frank. ( Thank you !)
By using network analyzer AG4395A with the impedance test kit AG43961A (these are at the PSL lab.), the impedance of resonant circuit with EOM was measured.
The picture of setup is attached. This impedance test kit allows to measure typically 0.1 [Ohm]-1M [Ohm] and frequency range of 100kHz-500MHz.
(result)
The resultant plot is attached. In the plot the blue curve represents the impedance measured by usual analyzer at 40m.
Note this curve is already subtracted the effect of the parasitic resistance.
( the parasitic resistance is in parallel to the circuit and it has ~7.8k Ohm, which is measured while the probe of the analyzer stays open. )
The red curve is the re-measured data using the impedance test kit.
The important point is that; these two peak values at the resonance around 40MHz show good agreement in 10%.
The resonant frequencies for two data differs a little bit, which might be the effect of a stray capacitance ( ~several [pF] )
The red curve has a structure around 80MHz, I think this comes from the non-coaxial cables, which connect the circuit and analyzing kit.
You can see these cables colored black and red in the picture.
( conclusion )
Our measurement with the subtraction of the parasitic resistance effect is working reliably ! |
Attachment 1: DSCN0421.JPG
|
|
Attachment 2: EOM.png
|
|
2449
|
Wed Dec 23 17:33:14 2009 |
rana | Update | ASS | OAF Feedaround ON and doing something good |
The Rai box ran out of batteries a couple of days ago and so the data is no good. I've put the Ranger back on the SR560 for now (but with the damping resistor removed, so the gain is 2x more than before). |
2448
|
Wed Dec 23 16:34:25 2009 |
Koji | Update | IOO | MCT QPD/MC REFL QPD disabled |
For a certain investigation of the sum/diff module for MCT QPD/MC REFL QPD, I removed it from the system.
|
2447
|
Tue Dec 22 18:42:40 2009 |
Sanjit, Koji | Configuration | Adaptive Filtering | Readded DAQ channels to active list |
Sometimes back we modified /cvs/cds/caltech/chans/daq/C1ASS.ini to save some of the channels. The file was reverted to default after the recent changes in ASS.
We again uncommented and made acquire=1 to save the following three channels using daqconfig:
C1:ASS-TOP_ERR_MCL_IN1_2048
C1:ASS-TOP_PEM_15_IN1_2048
C1:ASS-TOP_PEM_18_IN1_2048
The script automatically created a back up in /cvs/cds/caltech/chans/daq/archive
|
2446
|
Tue Dec 22 15:49:31 2009 |
kiwamu | Update | General | e-log restarted |
I found the e-log has been down around 3:40pm, then I restarted the e-log. Now it's working.
Thanks. |
2445
|
Tue Dec 22 13:14:37 2009 |
Alan | Omnistructure | PEM | miracle in the 40m |
Quote: |
We have been waiting for this for 20 some years. Arrowhead water with cooler. AWESOME
THANKS Alan
|
Happy holidays, everyone!
|
2444
|
Tue Dec 22 11:23:51 2009 |
kiwamu, Steve | Update | IOO | MC relocked |
In this morning I found MC unlocked.
Steve restored the watchdogs before I found that.
Then I relocked MC and now MC is locked and working well.
The reflected DC power is ~0.38, which is usual number.
|
2443
|
Tue Dec 22 10:04:41 2009 |
steve | Omnistructure | PEM | miracle in the 40m |
We have been waiting for this for 20 some years. Arrowhead water with cooler. AWESOME
THANKS Alan |
Attachment 1: h2oo.JPG
|
|
2442
|
Tue Dec 22 02:50:09 2009 |
rana, kiwamu, jenne | Update | ASS | OAF Feedaround ON and doing something good |
Kiwamu made the OAF screen functional today - screenshot attached.
After this, I used the measured TF of the MC1 to MCL to filter the signals from the Wilcoxon accelerometers and feed them into the MC.
The noise at 3 Hz went down by a factor of ~3. There's a little excess created at 100 Hz. Its good to see that our intuition about feed-forward is OK.
I did all of the filter calculations by adapting the scripts that Haixing, Valera, and I got going at LLO. They're all in the mDV/extra/C1 SVN.
The Wiener code predicts much better performance from using more than just 2 horizontal accelerometers, but I was too lazy to do more channels today.
I also added the Rai box to the Ranger readout today - the noise at 0.1 Hz went down by a factor of 10 and the noise at 1 Hz is close to 10^-11 m/rHz. |
Attachment 1: oaf.png
|
|
Attachment 2: oaf.png
|
|
2441
|
Mon Dec 21 19:24:29 2009 |
rana | Update | ASS | OAF Model update and build instructions |
I fit the MC1 -> MCL TF using vectfit4.m (from mDV). The wrapper file is mDV/extra/C1/ fitMC12MCL.m .
Plotted here are the data (RED), the fit (BLUE), and the residual x10 (GREEN).
For the magnitude plot, residual is defined as ------ res = 1 - fit / data
For the phase plot the residual is defined as ------- res = phase(data)-phase(fit)
You can see that the agreement is very good. The phase match is better than 5 deg everywhere below 10 Hz.
This TF is so smooth that we could have probably done without using this, but its good to excercise the method anyway. |
Attachment 1: mc12mcl.png
|
|
2440
|
Mon Dec 21 10:09:06 2009 |
jenne | Update | ASS | OAF Model update and build instructions |
OAF stands for Online Adaptive Filtering. We use the same computer which was once the ASS. One of these days, we'd like to completely be rid of all things which refer to ASS, and make even the computer's name OAF. |
2439
|
Mon Dec 21 08:16:54 2009 |
steve | Metaphysics | PEM | jackhammering next door |
At ITMX, on the CES side, 5 Ft from the wall the jackhammer is on. The susses are holding well. |
2438
|
Mon Dec 21 07:30:58 2009 |
??? | Update | ASS | OAF Model update and build instructions |
What does OAF stand for? The entry doesn't say that. Also the acronym is not in the abbreviation page of the wiki.
Can anyone please explain that? |
2437
|
Mon Dec 21 02:22:31 2009 |
rana, jenne, kiwamu | Update | ASS | OAF Model update and build instructions |
This allowed measuring the MC1 -> MCL TF finally. Its mostly flat. Data saved as Templates/OAF/OAF-MCLTF.xml |
Attachment 1: Untitled.png
|
|
2436
|
Mon Dec 21 01:14:08 2009 |
rana | Summary | Electronics | Noise measurement of the Rai Weiss FET preamp box |
I shorted the input to the box and then put its output into the SR560 (low noise, G = 100, AC). I put the output of the SR560 into the SR785.
*** BTW, the 2nd channel of the SR785 is kind of broken. Its too noisy by a factor of 100. Needs to go back for repair once we get started in the vac.
The attached PNG shows its input-referred noise with the short.
The picture shows the inside of the box before I did anything. The TO-5 package metal can is the meaty super dual-FET that gives this thing all of its low noise power.
 
In the spectra on the right are two traces. The BLUE one is the noise of the box as I found it. The BLACK one is the noise after I replaced R1, R6, R7, & R10 with metal film resistors.
The offset at the output of the box with either an open or shorted input is +265 mV.
I think we probably should also replace R2, R3, & R1, but we don't have any metal film resistors lower than 100 Ohms in the kit...but hopefully Steve will read this elog and do the right thing. |
Attachment 1: IMG_0242.JPG
|
|
2435
|
Sun Dec 20 23:42:44 2009 |
Jenne | Update | IOO | New Input Mode Matching Telescope |
I've got most of the new Mode Matching Telescope figured out. The scripts and an example result are at: MMT09 wiki (Rather, the scripts are in the svn: MMT svn)
Issues still to be resolved:
* We're getting pretty iffy 'angles' between tilt and translation when using the mode matching mirrors for steering.
* I haven't taken into account the astigmatism which occurs when you tilt the mode matching mirrors.
The nifty thing about these scripts is that they take a look at the mode matching overlap: For each possible mode matching solution it adds noise to all of the distances and radii of curvature during ~10,000 iterations and plots a histogram of the overlap so that we can see which solutions have a better chance of giving us the optimal overlap, even if we place the optics in slightly the wrong place.
I'd like to update the overlap part of the script with the astigmatism business: do we lose goodness of overlap if we tilt the mirrors by a bit? I think this will require redoing the overlap part with the X and Y directions separate. Koji has done this in the past. My current code assumes that the beam is always symmetric in X and Y. |
2434
|
Sun Dec 20 21:39:40 2009 |
rana, jenne, kiwamu | Update | ASS | OAF Model update and build instructions |
After a lot of headache, I got the OAF working again - read on for details.....................
Sometime last week, Jenne, Kiwamu, and I tried to update the OAF model to include the IIR "feed-around" path.
This path is in parallel to the existing FIR-based adaptive FXLMS stuff that Matt put in earlier. The reason for the
new path is that we want to try emulating the same FF technology which has been successful lately at LLO.
However, we were unable to make the ASS work after this work. Mostly, the build stuff worked fine, but we couldn't get DTT
to make a transfer function. The excitation channels could be selected and the excitation would actually start and get all the
way into MC1, but DTT would just hang on the first swep-sine measurement with no time-out error. Clearly our ASS building
documentation is no good. We tried using the instructions that Koji gave us for AAA, but that didn't completely work.
In particular, the 'make-uninstall-daq-ass' command gave this command:
[controls@c1ass advLigo]$ make uninstall-daq-ass
grep: target/assepics/assepics*.cmd: No such file or directory
Please make ass first
make: *** [uninstall-daq-ass] Error 1
re-arranging the order to do 'make-ass' first fixes this issue and so I have fixed this in the OAF Wiki.
The there's the whole issue with the tpchn_C3.par file. This contains all the test point definitions for the ASS/OAF machine. The main
IFO numbers are all in tpchn_C1.par and the OMC is all in tpchn_C2.par. When we do the usual build, in the 'make install-daq-ass' part:
[controls@c1ass advLigo]$ make install-daq-ass
Installing GDS node 3 configuration file
/cvs/cds/caltech/target/gds/param/tpchn_C3.par
Updating DAQ configuration file
/cvs/cds/caltech/chans/daq/C1ASS.ini
we get this .par file installed in the target area. The ACTUAL param file seems to actually be in /cvs/cds/caltech/gds/param !!
Of course, it still doesn't work. That's because the standard build likes to point to /cvs/cds/caltech/gds/bin/awgtpman and the one that runs on
linux is actually /opt/gds/awgtpman. So I've now made a file called startup_ass.cmd.good which runs the correct one. However, the default build
will try to start the wrong one and we have to fix the 'startass' script to point to the correct one on each build. Running the correct awgtpman
allows us to get the TP data using tools like tdsdata, so far no luck with DTT.\
UPDATE (23:33): It turns out that it was my old nemesis, NTPD. c1ass had a /etc/ntp.conf file that was pointing at an ntp server called rana113. I
am not an NTP server; I don't even know what time it is. I have fixed the ntp.conf file by making it the same ass c1omc (it now points to nodus). After
this I set the date and time manually (sudo date -s "20 DEC 2009 23:27:45" ) and then restarted NTPD. It should now be fine even when
we reboot c1ass.
After all of this nonsense, I am able to get TP data from c1ass and take transfer functions between it and the rest of the world ! |
2433
|
Sun Dec 20 14:34:24 2009 |
Koji | Update | SUS | ETMY watchdog tripped Sunday 5:00AM local |
It seemed that the ETMY watchdog tripped early Sunday morning.
The reason is not known. I just looked at ETMX, but it seemed fine.
I called the control room just in case someone is working on the IFO.
Also I did not see any elog entry to indicate on going work there.
So, I decided to reset the watchdog for ETMY. And it is working fine again. |
Attachment 1: Y.png
|
|
2432
|
Sat Dec 19 14:33:25 2009 |
Koji | Configuration | Computers | PDFlib lite / gnuplot 4.2.6 on Rosalba/Allegra |
In order to enable 'set terminal pdf' in gnuplot on Rosalba/Allegra, I installed PDFlib lite and gnuplot v4.2.6. to them.
(PDFlib lite is required to build the pdf-available version of gnuplot)
Installation of PDFlib lite:
- Building has been done at rosalba
- Download the latest distribution of PDFlib lite from http://www.pdflib.com/products/pdflib-7-family/pdflib-lite/
- Expand the archive. Go into the expanded directory
tar zxvf PDFlib-Lite-7.0.4p4.tar.gz
cd ./PDFlib-Lite-7.0.4p4
- configure & make
./configure
make
- install the files to the system / configure the dinamic linker
sudo make install
sudo ldconfig
Installation of gnuplot:
- Building has been done at rosalba
- Download the latest distribution of gnuplot form http://www.gnuplot.info/
- Expand the archive. Go into the expanded directory
tar zxvf gnuplot-4.2.6.tar.gz
cd ./ gnuplot-4.2.6.tar.gz
- configure & make
./configure --prefix=/cvs/cds/caltech/apps/linux/gnuplot
make
make install
- Create symbolic links of the executable at
/cvs/cds/caltech/apps/linux/bin
/cvs/cds/caltech/apps/linux64/bin
- Note: Although the original (non-PDF) gnuplot is still at
/usr/bin/gnuplot
new one is active because of the path setting
rosalba:linux>which gnuplot
/cvs/cds/caltech/apps/linux64/bin/gnuplot
|
2431
|
Fri Dec 18 15:40:33 2009 |
Koji | Update | IOO | MC2 spot centered / MCT QPD issue |
This afternoon I felt like saying hello to the input mode cleaner. So I decided to center the spot on MC2.
Motivation
MC has 6 alignment dofs. 4 of them are controlled by the WFSs. Remaining 2 appears at the spot position on MC2.
If the spot on the MC2 is fixed, the beam hits the same places of three mirrors. If the mirrors are completely fixed
in terms of the incident beam, I suppose the reflected beam is also fixed. This makes the WFS spots more stable.
Then I feel better.
Today's goal is to confirm the behaviour of MC such as dithering amplitude, response of the couplings to the alignment,
behavior of the WFS, and the transmitted power.
Method
1) Turned off MC auto locker. Turned off MC WFS as the WFS servos disturbs my work.
2) Dithered MC2 in Pitch and Yaw using DTT. There looks elliptic filter (fc=28Hz) in the ASC path, I used 20Hz-ish excitations.
- C1:SUS-MC2_ASCPIT_EXC 100cnt_pk@19Hz
- C1:SUS-MC2_ASCYAW_EXC 100cnt_pk@22Hz
3) Looked at C1:SUS-MC2_MCL_OUT to find the peaks at 19Hz and 22Hz. These are caused by alignment-length coupling.
If they are minimized I assume the spot is somehow centered on MC2.
Note: This may not be the true center. The suspension response should be investigated. But this is a certain reporoducible spot position.
Note: I should use ezcademod in order to obtain the phase information of the dither result.
4) Move MC2 Pitch for certain amount (0.01cnt) by the alignment slider. Align MC1/MC3 to have max transmittion.
5) If the Pitch peak got lower, the direction of 4) was right. Go further.
5') If the Pitch peak got higher, the direction of 4) was wrong. Go the other direction.
6) Repeat 4)&5) for Yaw.
Result
After the adjustment, the couplings got lower about 10 times. (Sorry! The explanation is not so scientific.)
Next time I (or someone) should make a script to do it and evaluate the coupling by the estimated distance of the spot from the center of the mirror (the center of the rotation).
I have not seen visible change in the spectrum of C1:SUS-MC2_MLC_OUT.
MCT QPD issue
By the spot centering, I could expected to see some improvement of the transmittion. But in reality, there was no change.
In fact, the transmittion power was getting down for those weeks.
I checked WFS and MCT paths. Eventually I found that a couple of possible problems:
1) MCT Total output varies more than 10% depending on the spot position on MCT QPD.
2) Just before the QPD, there is a ND1 filter.
This may suggest that:
a) Four elemtns of the MCT QPD have different responses.
b) The ND filter is causing a fringe.
So far I aligned the ND filter to face the beam. The reflection from the filter was blocked at a farther place.
Still the output varies on the spot position. Probably I have to look at the QPD someday.
So far the spot on the QPD was defined so that I get the maximum output from the QPD. This is about 8.8.
As I touched the steering mirrors, the X and Y outputs of the QPD are no longer any reference.
For now, I closed the PSL table. The full IFO was aligned. |
2430
|
Thu Dec 17 23:27:23 2009 |
rana | Update | PEM | Ranger Noise: sim w. Rai FET box as readout |
I have started measuring the low frequency noise of the FET front end + LT1128 low noise preamp from Rai Weiss. It has a very low input current noise because its FET based, which is not surprising. It is also a fairly low voltage noise box - the best measured ones have an input referred noise of ~0.35 nV/rHz.
Today I measured the noise of the one we have down to 0.1 Hz. It looks like a good candidate for a Geophone readout (e.g. Ranger or GS-13 or perhaps the L-4C). Because I didn't thermally shield any of this stuff, the broadband noise is ~0.8 nV/rHz. The low frequency corner is ~15 Hz.
I attach the LISO simulation of the voltage noise referred to the input. The circuit is described in this entry.
We can probably do better than this if we package it a little better or give it time to warm up or use metal film resistors inside. Even as it is, however, it would allow us to reach the thermal noise of the Ranger (or GS-13) down to 0.1 Hz.
This should be ~1.5 or 2x better than the LT1012 based readout at 1 Hz and 10x better down at 0.1 Hz (c.f. T0900457). |
Attachment 1: ranger.pdf
|
|
2429
|
Thu Dec 17 19:03:14 2009 |
Alberto | HowTo | Computers | Nodus sluggish |
Quote: |
The elog has been quite slow for the last two days. The cause is nodus, that has been slowing down the access to it.
I looked at the list of the running processes on nodus by typing the command prstat, which is the equivalent for Solaris of the Linux "top". I didn't see any particular process that might be absorbing too many resources.
I remember Rana fixing the same problem in the past but couldn't find his elog entry about that. Maybe we should just restart nodus, unless someone has some other suggestion.
|
Problem solved. Nodus and the elog are running fine. It's just that the elog takes some time to make a preview of complex pdf attachments, like those in Kiwamu's entry 2405. |
2428
|
Thu Dec 17 17:13:50 2009 |
Alberto | Omnistructure | Environment | STACIS stuff |
One of the electronics benches is currently occupied by the STACIS equipment.
We need that table If no one is working on the STACIS anymore, it should be removed from there. |
2427
|
Thu Dec 17 09:30:08 2009 |
Alberto | HowTo | Computers | Nodus sluggish |
The elog has been quite slow for the last two days. The cause is nodus, that has been slowing down the access to it.
I looked at the list of the running processes on nodus by typing the command prstat, which is the equivalent for Solaris of the Linux "top". I didn't see any particular process that might be absorbing too many resources.
I remember Rana fixing the same problem in the past but couldn't find his elog entry about that. Maybe we should just restart nodus, unless someone has some other suggestion. |